# ASAP: Advancing Semantic Alignment Promotes Multi-Modal Manipulation Detecting and Grounding

### v1

@@ 1 Introduction
A notable truth for the DGM4 task is that if there is manipulation in a piece of media, the image and text are usually not aligned any more, and this statement can be validated by the setup of this task [21]. With this critical observation, we argue that one of the primary keys for DGM4 models is to effectively perceive the fine-grained image-text alignment in the media. As depicted in Figure 1, if the model can capture that the word ”celebrating” does not align with the ”disappointment” emotion on the human face, then the model can acquire the valuable priori knowledge that the media may be manipulated. Existing methods predominantly rely on annotations to directly pinpoint the manipulated components, but they often simply apply the alignment learning on the unchanged multimodal media, which is not sufficiently effective to assist the subsequent detecting and grounding.
Addressing the aforementioned issues, this paper centers on improving fine-grained alignment learning for the DGM4 task. Two key challenges arise: firstly, manipulated media’s image and text are unpaired, precluding the use of conventional cross-modal alignment techniques such as contrastive learning. In response, we propose leveraging off-the-shelf Multimodal Large Language Models (MLLMs) to generate descriptions for both manipulated and non-manipulated images, followed by implementing contrastive learning between the generated captions and the respective images to fortify cross-modal alignment.
@@ 2 Relate Work
Large Models. Large models have made significant strides in natural language processing (NLP) and multimodal tasks. Large language models (LLMs) like Mistral [7] and LLaMA [24] demonstrate impressive performance, with Mistral excelling in language understanding and generation, and LLaMA achieving high efficiency with fewer parameters. In the multimodal domain, models like BLIP2 [9] excel in image understanding and generation, while multimodal GPT-4 [16] integrates text and image processing for enhanced capabilities across tasks. Google’s DeepMind Gato model [19] further pushes the boundaries with its general-purpose multimodal abilities. The growing adoption of large-scale models is driving innovation and providing novel solutions across various fields.
In this study, we leverage both LLMs and MLLMs to generate auxiliary texts, enhancing our model’s performance in multimodal tasks.
@@ 3 Methodology
{$&p(I,E)=\text{sigmoid}(\text{MLP}({\delta}F_{mm}^{v}+F_{mm}^{t})),\\ &F_{mm}^{v}=\mathcal{E}_{mm}^{v}(\mathcal{E}_{v}(I),\mathcal{E}_{t}(E)),\\ &F_{mm}^{t}=\mathcal{E}_{mm}^{t}(\mathcal{E}_{t}(E),\mathcal{E}_{v}(I)),\end{aligned}\right.$ (4)
@@ 3.3 Training
${P}_{i}=\begin{cases}1,&\text{if }I_{i}\text{ is manipulated},\\ 0,&\text{if }I_{i}\text{ is adjacent to manipulated patches},\\ -1,&\text{otherwise (would be ignored)}.\end{cases}$ (10)
@@
Tasks Reference Binary Cls Multi-Label Image Grounding Text Grounding $\Delta$AVG
@@
HAMMER++ [22] TPAMI24 93.33 14.06 86.66 86.41 79.73 80.71 76.46 83.77 76.03 73.05 72.14 72.59 +0.42
VIKI [11] IF24 93.51 13.87 86.67 86.58 81.07 80.10 76.51 83.95 75.77 77.79 66.06 72.44 +0.57
UFAFormer [13] IJCV24 93.81 13.60 86.80 87.85 80.31 81.48 78.33 85.39 79.20 73.35 70.73 72.02 +1.12
ASAP(Ours) - 94.38 12.73 87.71 88.53 81.72 82.89 77.35 84.75 76.54 79.38 73.86 76.52 +2.54
Table 1: Quantitative comparison of our proposed ASAP method with the current state-of-the-art techniques across four key tasks: manipulation detection, manipulation type identification, image grounding, and text grounding. As evident from the comparison, ASAP achieves the top performance, showing a significant average improvement over the baseline HAMMER method in all sub-tasks. $\Delta$AVG indicates the improvement compared to HAMMER.
@@ 6.2 Binary Classification
$&\mathcal{L}_{\text{IMG}}=\mathbb{E}_{(I,T)}\left[\boldsymbol{H}(C_{b}(M_{it}),y_{bin})\right]\\ &M_{it}={\delta}\mathcal{E}_{mm}^{v}(\mathcal{E}_{v}(I),\mathcal{E}_{t}(T))+\mathcal{E}_{mm}^{t}(\mathcal{E}_{t}(T),\mathcal{E}_{v}(I))\end{aligned}\right.$
@@ 6.3 Manipulation Type Detection
For the binary classification task, we input the multimodal feature $M_{it}$ into the Binary Classifier $C_{b}$ and compute the Binary Classifier Loss as:
@@ 6.4 Manipulated Text Token Grounding
$&\mathcal{L}_{\text{TMG}}=(1-\alpha)\mathcal{L}_{\text{tok}}+\alpha\mathcal{L}_{tok}^{mom}\\ &\mathcal{L}_{\text{tok}}=\mathbb{E}_{(I,T)}\left[\boldsymbol{H}(D_{t}(t_{tok}^{t}),y_{tok})\right]\\ &\mathcal{L}_{tok}^{mom}=\mathbb{E}_{(I,T)}KL\left[D_{t}(t_{tok}^{t})||\hat{D_{t}}(\hat{t}_{tok}^{t})\right]\\ &\{t_{cls}^{t},t_{tok}^{t}\}=\mathcal{E}_{mm}^{t}(\mathcal{E}_{t}(T),\mathcal{E}_{v}(I))\end{aligned}\right.$
@@ 7.1 Effectiveness of Cross-modality learning
Tasks Binary Cls Image Grounding

### v2

@@ 1 Introduction
A notable truth for the DGM4 task is that if there is manipulation in a piece of media, the image and text are usually not aligned any more, and this statement can be validated by the setup of this task [21]. With this critical observation, we argue that one of the primary keys for DGM4 models is to effectively perceive the fine-grained image-text alignment in the media [pfan, sigir24, shuyu]. As depicted in Figure 1, if the model can capture that the word ”celebrating” does not align with the ”disappointment” emotion on the human face, then the model can acquire the valuable priori knowledge that the media may be manipulated. Existing methods predominantly rely on annotations to directly pinpoint the manipulated components, but they often simply apply the alignment learning on the unchanged multimodal media, which is not sufficiently effective to assist the subsequent detecting and grounding.
Addressing the aforementioned issues, this paper centers on improving fine-grained alignment learning for the DGM4 task. Two key challenges arise: firstly, manipulated media’s image and text are unpaired, precluding the use of conventional cross-modal alignment techniques such as contrastive learning [pfan, pfan1, sigir24]. In response, we propose leveraging off-the-shelf Multimodal Large Language Models (MLLMs) to generate descriptions for both manipulated and non-manipulated images [vlp], followed by implementing contrastive learning between the generated captions and the respective images to fortify cross-modal alignment.
@@ 2 Relate Work
Large Models. Large models have made significant strides in natural language processing (NLP) and multimodal tasks. Large language models (LLMs) like Mistral [7] and LLaMA [24] demonstrate impressive performance, with Mistral excelling in language understanding and generation, and LLaMA achieving high efficiency with fewer parameters. In the multimodal domain, models like BLIP2 [9] excel in image understanding and generation, while multimodal GPT-4 [16] integrates text and image processing for enhanced capabilities across tasks. Google’s DeepMind Gato model [19] further pushes the boundaries with its general-purpose multimodal abilities. The growing adoption of large-scale models is driving innovation and providing novel solutions across various fields. In this study, we leverage both LLMs and MLLMs to generate auxiliary texts, enhancing our model’s performance in multimodal tasks.
@@ 3 Methodology
$\left\{\begin{aligned} &p(I,E)=\text{sigmoid}(\text{MLP}({\delta}F_{mm}^{v}+F_{mm}^{t})),\\ &F_{mm}^{v}=\mathcal{E}_{mm}^{v}(\mathcal{E}_{v}(I),\mathcal{E}_{t}(E)),\\ &F_{mm}^{t}=\mathcal{E}_{mm}^{t}(\mathcal{E}_{t}(E),\mathcal{E}_{v}(I)),\end{aligned}\right$. (4)
@@ 3.3 Training
[…]
@@
[…]
@@
HAMMER++ [22] TPAMI24 93.33 14.06 86.66 86.41 79.73 80.71 76.46 83.77 76.03 73.05 72.14 72.59 +0.39
VIKI [11] IF24 93.51 13.87 86.67 86.58 81.07 80.10 76.51 83.95 75.77 77.79 66.06 72.44 +0.37
UFAFormer [13] IJCV24 93.81 13.60 86.80 87.85 80.31 81.48 78.33 85.39 79.20 73.35 70.73 72.02 +1.13
ASAP(Ours) CVPR25 94.38 12.73 87.71 88.53 81.72 82.89 77.35 84.75 76.54 79.38 73.86 76.52 +2.40
[…]
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Key R&D Program of China (2024YFB3311602), the Fundamental Research Funds for the Central Universities, including JZ2024HGTB0261, JZ2023YQTD0072, JZ2024HGTG0309, and JZ2024AHST0337, the National Natural Science Foundation of China (72188101, 62020106007, 62302140), the Major Project of Anhui Province (202203a05020011), and the Anhui Provincial Natural Science Foundation - Distinguished Young Scholars Program (2408085J040). This work also received support from the New Cornerstone Science Foundation through the XPLORER PRIZE.
@@ 6.2 Binary Classification
$\left\{\begin{aligned} &\mathcal{L}_{\text{IMG}}=\mathbb{E}_{(I,T)}\left[\boldsymbol{H}(C_{b}(M_{it}),y_{bin})\right]\\ &M_{it}={\delta}\mathcal{E}_{mm}^{v}(\mathcal{E}_{v}(I),\mathcal{E}_{t}(T))+\mathcal{E}_{mm}^{t}(\mathcal{E}_{t}(T),\mathcal{E}_{v}(I))\end{aligned}\right$.
@@ 6.3 Manipulation Type Detection
For detection of the type of manipulation task, we input the multimodal feature $M_{it}$ into the Binary Classifier $C_{m}$ and compute the loss of the binary classifier as:
@@ 6.4 Manipulated Text Token Grounding
$\left\{\begin{aligned} &\mathcal{L}_{\text{TMG}}=(1-\alpha)\mathcal{L}_{\text{tok}}+\alpha\mathcal{L}_{tok}^{mom}\\ &\mathcal{L}_{\text{tok}}=\mathbb{E}_{(I,T)}\left[\boldsymbol{H}(D_{t}(t_{tok}^{t}),y_{tok})\right]\\ &\mathcal{L}_{tok}^{mom}=\mathbb{E}_{(I,T)}KL\left[D_{t}(t_{tok}^{t})||\hat{D_{t}}(\hat{t}_{tok}^{t})\right]\\ &\{t_{cls}^{t},t_{tok}^{t}\}=\mathcal{E}_{mm}^{t}(\mathcal{E}_{t}(T),\mathcal{E}_{v}(I))\end{aligned}\right$.
@@ 7.1 Effectiveness of Cross-modality learning
Tasks Binary Cls Text Grounding